OpenStudy (anonymous):

PLEASE HELP! METAL. Find a cubic function with the given zeros. Square root of seven., -Square root of seven., -4 f(x) = x3 + 4x2+ 7x - 28 f(x) = x3 + 4x2 - 7x - 28 f(x) = x3 - 4x2 - 7x - 28 f(x) = x3 + 4x2 - 7x + 28

OpenStudy (anonymous):

\[(x+\sqrt7)(x-\sqrt7)=x^2-7\]

OpenStudy (anonymous):

\[\sqrt{7},-\sqrt{7},- 4\]

OpenStudy (anonymous):

multiply that happiness by \((x+4)\) and you will have your answer

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i.e. \[(x+\sqrt7)(x-\sqrt7)(x+4)\] has the zeros you want then \[(x^2-7)(x+4)\] is all that is left to do

OpenStudy (anonymous):

if you are nervous about not getting the right answer use technology but it should not be too hard to do that multiplication

OpenStudy (anonymous):

f(x) = x3 + 4x2 - 7x - 28?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Lol, thanks.

OpenStudy (anonymous):

yw say hi to johnny for me

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

how do you know johnny, that's my BF.